# Tony Lindeberg

### v5

@@ 1 Introduction
Based on these theoretical results, we will then propose an extended set of theoretical predictions in Section 6, and relate them to the neurophysiological results, that we currently find as available, and also propose new neurophysiological and psychophysical experiments to address the questions posed from this theoretical study, and which previous neurophysiological and psychophysical experiments may not have been designed to address.
@@ 2 Relations to previous work
This approach does specifically have structural similarities to the recently developed area of geometric deep learning (Bronstein et al. 2021, Gerken et al. 2023), where deep networks are formulated from the constraint that they should be well-behaved under the influence of geometric image transformations. For examples of deep networks that are covariant under spatial scaling transformations, see Worrall and Welling (2019), Bekkers (2020), Sosnovik et al. (2020, 2021), Zhu et al. (2022), Jansson and Lindeberg (2022), Lindeberg (2022), Zhan et al. (2022), Wimmer et al. (2023) and Perzanowski and Lindeberg (2025).
@@ 4.3 Idealized spatial or spatio-temporal models for simple cells in the primary visual cortex
This form of the temporal smoothing function corresponds to using an infinite set of first-order integrators that are coupled in cascade, with the time constants chosen so as to specifically enable temporal scale covariance. The distribution parameter $c>1$, often chosen as $c=\sqrt{2}$ or $c=2$, in this temporal smoothing function reflects the ratio between adjacent discrete temporal scale levels in the corresponding temporal scale-space model. (For practical implementation purposes, the infinite number of first-order integrators can often be truncated after the 4 to 8 slowest time constants.)
@@
Let us initially disregard the effects of the spatial and the temporal derivative operators in the idealized spatio-temporal receptive field model (18) by setting the differentiation orders to $m=0$ and $n=0$, leading to the following form for the spatio-temporal smoothing kernel [^6] In this treatment, ${\mathbb{S}}_{+}^{2}$ denotes the set of symmetric positive definite $2\times 2$ matrices. $T\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\times{\mathbb{R}}_{+}\times{\mathbb{S}}_{+}^{2}\times{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ according to
@@ 5.3.4 Scale-normalized velocity-adapted temporal derivatives
    $\partial_{{\varphi^{\prime}}^{m},\scriptsize\mbox{norm}}\,\partial_{{t^{\prime}}^{n},\scriptsize\mbox{norm}}^{m}L^{\prime}(x^{\prime},t^{\prime};\;s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})=\\ =\partial_{\varphi^{m},\scriptsize\mbox{norm}}\,\partial_{t^{n},\scriptsize\mbox{norm}}^{m}L(x,t;\;s,\Sigma,\tau,v),$ (100)
@@
    $(\nabla_{x^{\prime},\scriptsize\mbox{affnorm}}\,\partial_{{t^{\prime}}^{n},\scriptsize\mbox{norm}}^{m}L^{\prime})(x^{\prime},t^{\prime};\;s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})=\\ =\tilde{\rho}\,(\nabla_{x,\scriptsize\mbox{affnorm}}\,\partial_{t^{n},\scriptsize\mbox{norm}}^{m}L)(x,t;\;s,\Sigma,\tau,v)$ (109)
@@
    $({\cal H}_{x^{\prime},\scriptsize\mbox{affnorm}}\,\partial_{{t^{\prime}}^{n},\scriptsize\mbox{norm}}^{m}L^{\prime})(x^{\prime},t^{\prime};\;s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})=\\ =\tilde{\rho}\,({\cal H}_{x,\scriptsize\mbox{affnorm}}\,\partial_{t^{n},\scriptsize\mbox{norm}}^{m}L)(x,t;\;s,\Sigma,\tau,v)\,\tilde{\rho}^{T},$ (110)
@@ 5.5 Explicit examples of covariant receptive field families
Notably, these theoretical results comprise combinations of spatial derivatives up to order 2 with temporal derivatives for any order of temporal differentiation. In these ways, we can thus formulate a rich set of both purely spatial and joint spatio-temporal receptive field models, that are provably covariant under combinations of 4 main types of primitive geometric image transformations according to Equations (1)–(4), as summarized in the composed geometric image transformations according to (5), (6) and (9).
@@ 5.6 Relationships to the influence of illumination variations
In this context, it is interesting to note that the retinex theory of early vision (Land 1974; 1986) also makes use of a logarithmic brightness scale. An exposure mechanism on the retina that adapts the diameter of the pupil and the sensitivity of the photopigments in such a way that the relative range in the variability of the signal divided by the mean illumination is held constant, can also be seen as implementing an approximation of the derivative of a logarithmic transformation
@@ 6.1 Purely spatial variabilities in the shapes of spatial and spatio-temporal receptive fields
From the structure of orientation maps in the primary visual cortex of higher mammals, as studied by Bonhoeffer and Grinvald (1991), Blasdel (1992), Koch et al. (2016) and others, it is clear that we can interpret these orientation maps as an expansion of the receptive field shapes over the image orientations, corresponding to the parameter $\varphi$ in the idealized receptive field models (18) and (4.3). Thereby, we could regard the vision system of higher mammals to have the ability to compute visual operations corresponding to rotation covariance.
@@ 6.2 Additional variabilities involving the temporal domain
Figure 6 in this paper shows an example of such a variability under Galilean transformations for spatio-temporal receptive fields over a 1+1-D spatio-temporal domain, based on a first-order spatial derivative of a Gaussian kernel and a first-order temporal derivative of the time-causal limit kernel. Figure 18 shows the variability in the directional selectivity properties of simple cells that such a variability in the image velocity $v$ of the receptive fields gives rise to, when also including a variability over the degree of elongation $\kappa$ of the receptive fields, as based on the theoretical analysis in Lindeberg (2025g).
@@
  • If there would be expansions over both the degrees of freedom of the affine group beyond uniform scaling transformations and rotations as well as an expansion over the degrees of freedom of the Galilean group, are those expansions then performed jointly or separately? In other words, are such expansions performed over the tensor product of the corresponding affine and Galilean parameters or separately for each subset of these degrees of freedom?
@@ 7 Summary and discussion
Hence, a vision system that is able to extract these parameters of the geometric image transformations based on processing and comparing populations of receptive field responses, should in principle have the ability to compute direct cues to the structure of environment, directly from established matching relations over the receptive field responses between the different views, or in relation to a learned memory of receptive field responses from previous views. Thereby, certain functionalities of the vision system could be formulated directly in terms of the parameters of the primitive geometric image transformations between different views of the same scene or the same spatio-temporal event.
Thus, in summary, we have presented a coherent theory for how the structure of the geometric image transformations, that determine the formation of the image data that reach the visual sensors in the perception system, influence the desired structure of the receptive field responses in an idealized vision system. Specifically, we have argued how we based on the requirement of formulating identity relations between receptive field responses acquired under different viewing conditions for the same scene, are lead to the requirement of covariance properties of the receptive fields, which in turn imply a set of variabilities of the shapes of the receptive fields, to span the degrees of freedom of the corresponding geometric image transformations. In this way, we have formulated a set of theoretical predictions, and also discussed partial support of these predictions in view of existing neurophysiological evidence and complementary theoretical analysis in terms of cascade smoothing properties.

### v6

@@ 1 Introduction
Based on these theoretical results, we will then propose an extended set of theoretical predictions in Section 6, and relate them to the neurophysiological results that we currently find as available, and also propose new neurophysiological and psychophysical experiments to address the questions posed from this theoretical study, and which previous neurophysiological and psychophysical experiments may not have been designed to address.
@@ 2 Relations to previous work
This approach does specifically have structural similarities to the recently developed area of geometric deep learning (Bronstein et al. 2021, Gerken et al. 2023), where deep networks are formulated from the constraint that they should be well-behaved under the influence of geometric image transformations. For examples of deep networks that are covariant (equivariant) under spatial scaling transformations, see Worrall and Welling (2019), Bekkers (2020), Sosnovik et al. (2020, 2021), Zhu et al. (2022), Jansson and Lindeberg (2022), Lindeberg (2022), Zhan et al. (2022), Wimmer et al. (2023) and Perzanowski and Lindeberg (2025).
@@ 4.3 Idealized spatial or spatio-temporal models for simple cells in the primary visual cortex
This form of the temporal smoothing function corresponds to using an infinite set of first-order integrators that are coupled in cascade, with the time constants chosen so as to specifically enable temporal scale covariance. The distribution parameter $c>1$, often chosen as $c=\sqrt{2}$ or $c=2$, in this temporal smoothing function reflects the ratio between adjacent discrete temporal scale levels in the corresponding temporal scale-space model. For practical implementation purposes, the infinite number of first-order integrators can often be truncated after the 4 to 8 slowest time constants.
@@
Let us initially disregard the effects of the spatial and the temporal derivative operators in the idealized spatio-temporal receptive field model (18), by setting the differentiation orders to $m=0$ and $n=0$, leading to the following form for the spatio-temporal smoothing kernel [^6] In this treatment, ${\mathbb{S}}_{+}^{2}$ denotes the set of symmetric positive definite $2\times 2$ matrices. $T\colon{\mathbb{R}}^{2}\times{\mathbb{R}}\times{\mathbb{R}}_{+}\times{\mathbb{S}}_{+}^{2}\times{\mathbb{R}}_{+}\times{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ according to
@@ 5.3.4 Scale-normalized velocity-adapted temporal derivatives
    $\partial_{{\varphi^{\prime}}^{m},\scriptsize\mbox{norm}}\,\partial_{\bar{t^{\prime}},\scriptsize\mbox{norm}}^{n}L^{\prime}(x^{\prime},t^{\prime};\;s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})=\\ =\partial_{\varphi^{m},\scriptsize\mbox{norm}}\,\partial_{\bar{t},\scriptsize\mbox{norm}}^{n}L(x,t;\;s,\Sigma,\tau,v),$ (100)
@@
    $(\nabla_{x^{\prime},\scriptsize\mbox{affnorm}}\,\partial_{\bar{t^{\prime}},\scriptsize\mbox{norm}}^{n}L^{\prime})(x^{\prime},t^{\prime};\;s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})=\\ =\tilde{\rho}\,(\nabla_{x,\scriptsize\mbox{affnorm}}\,\partial_{\bar{t},\scriptsize\mbox{norm}}^{n}L)(x,t;\;s,\Sigma,\tau,v)$ (109)
@@
    $({\cal H}_{x^{\prime},\scriptsize\mbox{affnorm}}\,\partial_{\bar{t^{\prime}},\scriptsize\mbox{norm}}^{n}L^{\prime})(x^{\prime},t^{\prime};\;s^{\prime},\Sigma^{\prime},\tau^{\prime},v^{\prime})=\\ =\tilde{\rho}\,({\cal H}_{x,\scriptsize\mbox{affnorm}}\,\partial_{\bar{t},\scriptsize\mbox{norm}}^{n}L)(x,t;\;s,\Sigma,\tau,v)\,\tilde{\rho}^{T},$ (110)
@@ 5.5 Explicit examples of covariant receptive field families
Notably, these theoretical results comprise combinations of spatial derivatives up to order 2 with temporal derivatives for any order of temporal differentiation (including the special case when the temporal order of differentiation $n=0$). In these ways, we can thus formulate a rich set of both purely spatial and joint spatio-temporal receptive field models, that are provably covariant under combinations of 4 main types of primitive geometric image transformations according to Equations (1)–(4), as summarized in the composed geometric image transformations according to (5), (6) and (9).
@@ 5.6 Relationships to the influence of illumination variations
In this context, it is interesting to note that the retinex theory of early vision (Land 1974; 1986) also makes use of a logarithmic brightness scale. An exposure mechanism on the retina that adapts the diameter of the pupil and the sensitivity of the photopigments, in such a way that the relative range in the variability of the signal divided by the mean illumination is held constant, can also be seen as implementing an approximation of the derivative of a logarithmic transformation
@@ 6.1 Purely spatial variabilities in the shapes of spatial and spatio-temporal receptive fields
From the structure of orientation maps in the primary visual cortex of higher mammals, as studied by Bonhoeffer and Grinvald (1991), Blasdel (1992), Koch et al. (2016) and others (see Figure 15 for an illustration), it is clear that we can interpret these orientation maps as an expansion of the receptive field shapes over the image orientations, corresponding to the parameter $\varphi$ in the idealized receptive field models (18) and (4.3). Thereby, we could regard the vision system of higher mammals to have the ability to compute visual operations corresponding to rotation covariance.
@@ 6.2 Additional variabilities involving the temporal domain
Figure 6 in this paper shows an example of such a variability under Galilean transformations for spatio-temporal receptive fields over a 1+1-D spatio-temporal domain, based on a first-order spatial derivative of a Gaussian kernel and a first-order temporal derivative of the time-causal limit kernel. Figure 18 shows the variability in the directional selectivity properties of simple cells, that such a variability in the image velocity $v$ of the receptive fields gives rise to, when also including a variability over the degree of elongation $\kappa$ of the receptive fields, as based on the theoretical analysis in Lindeberg (2025g).
@@
  • If there would be expansions over both the degrees of freedom of the affine group beyond uniform scaling transformations and rotations as well as an expansion over the degrees of freedom of the Galilean group, are those expansions then performed jointly or separately? In other words, are such expansions performed over the tensor product of the corresponding affine and Galilean parameters, or separately for each subset of these degrees of freedom?
@@ 7 Summary and discussion
Hence, a vision system, that is able to extract these parameters of the geometric image transformations based on processing and comparing populations of receptive field responses, should in principle have the ability to compute direct cues to the structure of environment, directly from established matching relations over the receptive field responses between the different views, or in relation to a learned memory of receptive field responses from previous views. Thereby, certain functionalities of the vision system could be formulated directly in terms of the parameters of the primitive geometric image transformations between different views of the same scene or the same spatio-temporal event.
Thus, in summary, we have presented a coherent theory for how the structure of the geometric image transformations, that determine the formation of the image data that reach the visual sensors in the perception system, influence the desired structure of the receptive field responses in an idealized vision system. Specifically, we have argued how we, based on the requirement of formulating identity relations between receptive field responses acquired under different viewing conditions for the same scene, are lead to the requirement of covariance properties of the receptive fields, which in turn imply a set of variabilities of the shapes of the receptive fields, to span the degrees of freedom of the corresponding geometric image transformations. In this way, we have formulated a set of theoretical predictions, and also discussed partial support of these predictions in view of existing neurophysiological evidence and complementary theoretical analysis in terms of cascade smoothing properties.
@@
7.1 Perspective and outlook
Let us conclude by relating the presented theory to a set of other related theories and concepts regarding vision.
Towards his goal of formulating an computational theory of vision, Marr (1982) proposed that the vision system may be creating a symbolic representation of intensity changes in terms of a primal sketch. While Marr primarily developed his notions around edge, blob and bar primitives using the difference-of-Gaussians (approximating the Laplacian-of-the-Gaussian) as a spatial filter, and which mimics the spatial component in many receptive fields in the retina and the lateral geniculate nucleus (LGN), with the more developed knowledge of the receptive fields in the primary visual cortex that we have today, we can instead regard the idealized models of spatial and spatial receptive fields used in this treatment as providing a much richer set of basis functions for registering the intensity variations over space and time as the input for later stage visual processes in the visual pathway. Specifically, the theory presented in this treatment goes substantially further, by also addressing intimate connections between the receptive fields and 3-D properties of the environment, such as surface orientation, variations in depth and object motion, and notably incorporating the effects of the perspective and the projective image transformations directly into the filter shapes. Furthermore, although many of these receptive fields will often give rise to strong responses near edges, we may not regard their primary purpose as being edge detectors, but instead as members of a richer family of spatial and spatial receptive fields applicable for a multiple purposes in visual perception.
According to the sparse coding model for visual receptive fields (Olshausen and Field 1996; 1997), the receptive fields are suggested to be shaped like oriented edges partly for statistical efficiency, in order to compress natural images according to the redundancy reduction principle. This idea does also bear very close relationship to the efficient coding hypothesis for sensory neurons proposed by Barlow (1961). According to the here considered normative theory for visual receptive fields, the spatial primitives ought to, however, be rather shaped as affine Gaussians derivatives, to obey geometric covariance properties to handle variations in the viewing direction relative to 3-D objects.
According to the underlying normative theory for visual receptive fields, there is a set of idealized receptive field models to handle the influence of geometric image transformations. The underlying biological substrate in terms of simple cells may, on the other hand, have drifting properties over time. One may therefore ask if a biological vision system exposed to variabilities in image data, as generated from substantial variabilities in the underlying geometric image transformations, would converge to these idealized receptive field models by plasticity. In fact, for learning-based algorithms for computer vision, it has been demonstrated that idealized receptive field models structurally closely related to Gaussian derivatives have been demonstrated to well model the receptive fields learned by state-of-the-art convolutional networks when exposed the ImageNet dataset (Lindeberg et al. 2025), however, then with much lower variabilities in that dataset compared to the type of visual data seen by a possibly moving visual observer in a dynamic world.
Given the structurally very close connections between the predictions from the underlying normative theory for visual receptive fields, derived from structural (physical) properties of the environment, and properties of visual neurons in the primary visual cortex, one may hence consider if the vision systems of higher mammals have either implicitly or explicitly internalized structural/physical properties of the environment, either by evolution or as obtained by learning from the visual data that the vision system is exposed to from birth. Specifically, one may ask if we from this conceptual background can regard the influence of geometric image transformations as a primary factor for the development of the vision systems for higher mammals.
More generally, researchers have for decades worked on finding geometric invariants for visual perception, see e.g., Mundy and Zisserman (1992), Rolls (1994), Booth and Rolls (1998), Lowe (2004), Quiroga et al. (2005), DiCarlo and Cox (2007), Tuytelaars and Mikolajczyk (2008), Goris and Op de Beeck (2009), Lindeberg (2013b), Isik et al. (2013), Poggio and Anselmi (2016) and Rodríguez et al. (2018). From such a background, the covariant receptive field families considered in this work behave in a predictable way under the basic classes of geometric image transformations in vision, and do in this respect constitute a precursor for computing corresponding invariant visual representations at higher levels in the visual hierarchy.
Acknowledgements
I would like to thank Benjamin Auffarth for valuable suggestions regarding the conceptual desciption in Section 7.1.